OpenStudy (anonymous):

The surface area of a sphere is decreasing at the constant rate of 3π sq. cm/sec . At what rate is the volume of the sphere decreasing at the instant its radius is 2 cm ?

OpenStudy (anonymous):

getting stuck on this site not loading what is the surface area of a sphere?

OpenStudy (anonymous):

im searching on it

OpenStudy (anonymous):

oh doh it is the anti derivative of the volume making it \[S=4\pi r^2\]

OpenStudy (anonymous):

so basically i just substitute now?

OpenStudy (anonymous):

yeah since \[V'=4\pi r^2r'\]

OpenStudy (anonymous):

dont i have to solve for dr/dt?

OpenStudy (anonymous):

well let me think maybe not since the derivative of the volume is the surface are but maybe

OpenStudy (anonymous):

but you are told the rate of change of the surface area right? that is what is fixed so maybe it is a good idea to write the volume in terms of the surface area and skip the radius all together

OpenStudy (anonymous):

\[\large V=\frac{1}{6\sqrt{\pi}}S^{\frac{3}{2}}\] i think

OpenStudy (anonymous):

so it ends up like this?, 4pi(2)^(2) and what is r^I

OpenStudy (anonymous):

then \[V'=\frac{1}{4\sqrt{\pi}}\sqrt{S}S'\]

OpenStudy (anonymous):

the answer must be in cubic cm/sec

OpenStudy (anonymous):

you are given \(S'=-3\pi\) or something (hard for me to read) so it is easier so skip the radius in this

OpenStudy (anonymous):

ok so what would be the final answer?

OpenStudy (anonymous):

yes of course because it is a volume volume per second

OpenStudy (anonymous):

is the rate \(-3\pi\) ?

OpenStudy (anonymous):

\[S'=-3\pi\] \[r=2\] \[S=16\pi\] put them in here \[V'=\frac{1}{4\sqrt{\pi}}\sqrt{S}S'\]

OpenStudy (anonymous):

in fact now we see we didn't need any of this the derivative of the volume IS the surface area you get \(-3\pi\)

OpenStudy (anonymous):

i just typed it in my calculator and did the procedure like you just did thanks, im going to be posting some questions